# Geometric Freeze-Tag Problem

### v2

@@ Abstract.
We present new upper bounds for the $l_{1}$ and $l_{2}$ norms in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. For $(\mathbb{R}^{2},l_{2})$, we achieve a makespan of at most $5.064r$, improving on the previous bound of $7.07r$ by Bonichon et al. (DISC’24). In $(\mathbb{R}^{3},l_{1})$, we establish an upper bound of $13r$, which leads to a bound of $22.52r$ for $(\mathbb{R}^{3},l_{2})$. Here, $r$ denotes the maximum distance of a robot from the initially active robot under the given norm. To the best of our knowledge, these are the first known bounds for the makespan in $\mathbb{R}^{3}$ under these norms.
Additionally, we demonstrate that the maximum makespan for $n$ robots is not achieved when they are equally distributed on the boundary in $(\mathbb{R}^{2},l_{2})$. We also explore the FTP in $(\mathbb{R}^{3},l_{2})$ for specific instances where robots are positioned on a boundary, providing further insights into practical scenarios.
@@
Freeze-Tag Problem, Makespan, Approximation Algorithms, Robot Routing, Scheduling, Binary Trees, Distributed Computing
@@ 1. Introduction
Despite extensive research on the complexity of FTP, few results address the wake-up ratio. For $(\mathbb{R}^{2},l_{1})$, Bonichon et al. proved that the wake-up ratio is at most $5$ and provided an example with a makespan of $5$ Bonichon et al. (2024) (DISC’24). This gives an upper bound of $5\sqrt{2}$ for the wake-up ratio in $(\mathbb{R}^{2},l_{2})$, improving the previous upper bound of $10.06$ found in Yazdi et al. (2015).
@@ 2. Our Results
In this paper, we begin by focusing on FTP within the unit $l_{2}$-disk and introduce a new wake-up strategy. Our main result is an improved makespan, reducing the previous bound from $7.07$ to $5.064$, as established in Bonichon et al. (2024). Formally, we state the following theorem:
@@
A robot at the origin can wake up any set of $n$ asleep robots in the unit $l_{2}$-disk with a makespan of at most $5.064$. The proof of this theorem is provided in Section 3.1.
Next, we examine FTP in $\mathbb{R}^{3}$ and propose a new strategy for $(\mathbb{R}^{3},l_{1})$. We establish an upper bound of $13$ for $(\mathbb{R}^{3},l_{1})$, which leads to an upper bound of $13\sqrt{3}$ for $(\mathbb{R}^{3},l_{2})$. To our knowledge, no previous bounds for the makespan in $\mathbb{R}^{3}$ have been provided. Formally, we state the following theorem:
@@
A robot at the origin can wake up any set of $n$ asleep robots in the unit $l_{1}$-ball in $\mathbb{R}^{3}$ with a makespan of at most $13$. The proof of this theorem is presented in Section 4.
Our approaches for $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ are fundamentally different. In $\mathbb{R}^{2}$, we start at the origin and initially awaken robots within a disk of radius zero, since only the origin robot is awake. In each step, we expand the radius of this disk, waking up the robots that are closer to the origin, until the radius reaches $1$. By the end, all robots inside the disk are awake. However, this method cannot be extended to $\mathbb{R}^{3}$. Therefore, we employ a different strategy for $\mathbb{R}^{3}$.
@@
Next, we show that for certain values of $n$, the wake-up ratio for $n$ points is not achieved when the points are equally distributed on the unit circle. We do this by presenting specific instances and computing their makespan. This addresses a statement from Bonichon et al. (2024), which proposes: ”One of the following statements is surely wrong: The wake-up ratio for $n$ points is achieved by their equal distribution on the unit circle. For every $n\geq 4,\gamma_{n}(\mathbb{R}^{2},l_{2})>\gamma_{n+2}(\mathbb{R}^{2},l_{2})$.” We provide examples showing that the first statement is incorrect.
Finally, we examine a version of FTP in $(\mathbb{R}^{3},l_{2})$, where the asleep robots are located on the boundary of the unit $l_{2}$-ball. Using our approach for $(\mathbb{R}^{2},l_{2})$, we show that FTP in $(\mathbb{R}^{3},l_{2})$ for asleep robots on the boundary can be solved with a makespan of at most $9.26$. Formally, we have: {theorem} A robot at the origin can wake up any set of $n$ asleep robots on the boundary of the unit $l_{2}$-ball in $\mathbb{R}^{3}$ with a makespan of at most $9.26$. The proof of this theorem is deferred to Section 5.2.
Our approach involves mapping the points on the boundary of the unit $l_{2}$-ball in $\mathbb{R}^{3}$ to a $l_{2}$-disk in $\mathbb{R}^{2}$. By solving the problem in $\mathbb{R}^{2}$, we obtain a wake-up strategy for $\mathbb{R}^{3}$. One motivation for tackling the problem in this setting is our conjecture that the maximum makespan for $n$ points is reached when the points lie on the boundary of the unit $\eta$-ball.
A similar wake-up strategy allows us to tackle another variant of FTP, which we call surface-FTP. In this version, the points are positioned on the boundary of a unit $l_{2}$-ball in $\mathbb{R}^{3}$, with the active robot also on the surface. The distance between two points $v$ and $u$ is defined as the geodesic distance, or the length of the shortest arc connecting them on the surface. This problem has practical applications in areas like communication and transportation on Earth’s surface, where geodesic distance between points is important. Formally, we have: {theorem} Given an instance of surface-FTP, the makespan is at most $11.1$. This theorem is proven in Section 5.3.
@@ 3. FTP in $(\mathbb{R}^{2},l_{2})$
3.1. The wake-up ratio is at most $5.064$ in $(\mathbb{R}^{2},l_{2})$
This section establishes an upper bound of $5.064$ for $\gamma_{({\mathbb{R}^{2}},l_{2})}$. We present two key strategies for computing a wake-up tree and combine them to achieve an improved makespan.
@@ 3.1.1. Arc-Strategy
The Arc-Strategy provides an upper bound of $5.9851$ for the wake-up ratio in $(\mathbb{R}^{2},l_{2})$.
@@
$\sum_{i=1}^{k}{d_{i}}=d_{1}+\sum_{i=2}^{k}{\sqrt{b_{i}^{2}+a_{i}^{2}+2a_{i}b_{i}\sin{\frac{\alpha_{i-1}}{2}}}}$
@@
$\sum_{i=1}^{k}{d_{i}}$ $=d_{1}+$ $\sum_{i=2}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}(\sin{\frac{\alpha_{i-1}}{2}})^{2}}}$
@@
$\sum_{i=2}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}(\sin{\frac{\alpha_{i-1}}{2}})^{2}}}$
$\leq\sum_{i=2}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}}$
$=\sum_{i=2}^{k}{\sqrt{(a_{i}+2r_{i-1}\sin{\frac{\alpha_{i-1}}{2}})^{2}}}=\sum_{i=2}^{k}{a_{i}+2r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}$
@@
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $d_{1}+\sum_{i=2}^{k}{a_{i}}+2\sum_{i=2}^{k}{r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}$
Note that $\sum_{i=1}^{k}a_{i}=1$ and $a_{1}=d_{1}$ so by replacing $d_{1}$ and considering all $r_{i}\leq 1$ we have:
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\sum_{2=1}^{k}{\sin{\frac{\alpha_{i}}{2}}}+2r_{1}\sin{\frac{\alpha_{1}}{2}}$
Given that $\alpha_{i}\leq\frac{\pi}{2^{i-1}}$, we can write:
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\sum_{i=2}^{k}{\sin{\frac{\pi}{2^{i}}}}+2r_{1}\sin{\frac{\pi}{2}}$
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\left(r_{1}\sin{\frac{\pi}{2}}+\sin{\frac{\pi}{4}}+\dots\right)$
Using the fact that for $0\leq x\leq 1$, $\sin{x}\leq x$, we have:
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\left(r_{1}+\frac{\sqrt{2}}{2}\right)+\pi\left(\frac{1}{8}+\frac{1}{16}+\dots+\frac{1}{2^{n}}\right)$
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\left(r_{1}+\frac{\sqrt{2}}{2}\right)+\pi\left(\frac{1}{8}+\frac{1}{16}+\dots\right)$
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\left(r_{1}+\frac{\sqrt{2}}{2}+\frac{\pi}{4}\right)$
$\sum_{i=1}^{k}{d_{i}}$ $<$ $3.9851+2r_{1}\leq 5.9851$
$d_{1}$$r_{2}$$b_{2}$$a_{2}$$d_{2}$$\beta_{1}$$\beta_{1}$$\alpha_{1}$$p_{1}$$p_{2}$$p_{3}$$p_{4}$
@@
Thus, using the Arc-Strategy, FTP in $(\mathbb{R}^{2},l_{2})$ can be solved in at most $5.9851r$ time units, where $r$ is the radius of the disk containing all the asleep robots, and the active robot is at the center.
@@ 3.1.2. Ring-Strategy
  • Consider the projections of the points in $p_{0}$’s half of the ring onto the circle $C(O,r_{1})$. The projection of a point $p_{j}$ is the intersection of $\overline{Op_{j}}$ with the circle $C(O,r_{1})$. Let $p_{1}$ be the point whose projection is closest to $p_{0}$. Then, $p_{0}$ activates $p_{1}$.
@@
We now calculate the wake-up time for this strategy, specifically the time required to activate all the robots in a half-ring when the active robot is $p_{0}$. Since there are two awake robots at the same initial position, $p^{\prime}_{0}$ activates the robots in its own half in parallel. Therefore, the total wake-up time is the same for both halves, and all robots will be awake within this time.
@@ 3.1.3. Our combined strategy for FTP in $(\mathbb{R}^{2},l_{2})$
If $r_{1}\leq 0.539125$, we apply the Arc-Strategy, Lemma 3.1.1, resulting in a makespan of less than $3.9851+2r_{1}\leq 5.06335$. Since $r_{1}\leq 0.539125$, the desired makespan is achieved. If $r_{1}>0.539125$, the robot at the center, $p_{0}$, moves to the nearest robot, which takes $r_{1}$ time units. At this point, all remaining asleep robots are located within a ring with an inner radius of $r_{1}$ and an outer radius of $1$. By applying the Ring-Strategy (as detailed in Lemma 3.1.2), the robots can be activated in $\pi+3-3r_{1}$ time units. Therefore, the total time to activate all robots is at most $r_{1}+\pi+3-3r_{1}\leq 5.06335\simeq 5.064$.
@@ 3.2. Some examples
We start with the following statement of Bonichon et al. (2024) which says that one of the following statements are surely wrong:
(1) the wake-up ratio is attained for points that are equally distributed on unit circle.
(2) for every $n\geq 4$, $\gamma_{n}(l_{2})>\gamma_{n+1}(l_{2})$.
We present two instances that show the first statement is wrong. Our first instance is a set of $5$ asleep robots, the makespan of this instance is $3.530$, see Figure 5(a). When five points are equally distributed on the cycle then the makespan is $3.351$. Our second example is a set of $7$ asleep robots, the makespan of this instance is $3.498$, see Figure 5(b). When seven points are equally distributed on the cycle then the makespan is $3.431$. The makespans of these instances are computed using a computer program. The code calculates the makespan using brute force search. Our code is publicly available online [^1] Makespan Calculator Code.
$p_{0}$$p_{1}$$p_{2}$$p_{3}$$p_{4}$$p_{5}$
Point $\theta$ x y
$p_{1}$ $17.37$ $0.954$ $0.298$
$p_{2}$ $96.22$ $-0.108$ $0.994$
$p_{3}$ $174.42$ $-0.995$ $0.097$
$p_{4}$ $219.75$ $-0.768$ $-0.639$
$p_{5}$ $299.27$ $0.488$ $-0.872$
(a) Coordinates of five points with a makespan of $3.530$.
$p_{0}$$p_{1}$$p_{2}$$p_{3}$$p_{4}$$p_{5}$$p_{6}$$p_{7}$
@@
(b) Coordinates of seven points with a makespan of $3.498$.
@@
Our program computes the makespan for a given instance, and we’ve also searched for instances with large makespans. Furthermore, we generated over $10^{5}$ random instances for different values of $n$ and computed their makespans, leading to the following conjecture:

### v4

@@ Abstract.
We present new upper bounds for the $l_{1}$ and $l_{2}$ norms in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. For $(\mathbb{R}^{2},l_{2})$, we achieve a makespan of at most $5.4162r$, improving on the previous bound of $7.07r$ by Bonichon et al. (DISC’24). In $(\mathbb{R}^{3},l_{1})$, we establish an upper bound of $13r$, which leads to a bound of $22.52r$ for $(\mathbb{R}^{3},l_{2})$. Here, $r$ denotes the maximum distance of a robot from the initially active robot under the given norm. To the best of our knowledge, these are the first known bounds for the makespan in $\mathbb{R}^{3}$ under these norms.
We also explore the FTP in $(\mathbb{R}^{3},l_{2})$ for specific instances where robots are positioned on a boundary, providing further insights into practical scenarios.
@@
Freeze-Tag Problem; Makespan; Approximation Algorithms; Robot Routing; Scheduling; Binary Trees; Distributed Computing
@@ 1. Introduction
[…]
The unit ball in L1 and R3. Its shape looks like a prism, resembeling Ramiel from Neon Genesis Evangelion.
@@
[…]
An instance of FTP alongside it’s respective wake-up tree. On the right there is a 2D instance of FTP, containing the points in a unit disk, on the left we have a binary tree representing the wake-up tree for this scenario.
Despite extensive research on the complexity of FTP, few results address the wake-up ratio. For $(\mathbb{R}^{2},l_{1})$, Bonichon et al. proved that the wake-up ratio is at most $5$ and provided an example with a makespan of $5$ Bonichon et al. (2024a). This gives an upper bound of $5\sqrt{2}$ for the wake-up ratio in $(\mathbb{R}^{2},l_{2})$, improving the previous upper bound of $10.06$ found in Yazdi et al. (2015).
@@ 2. Our Results
In this paper, we begin by focusing on FTP within the unit $l_{2}$-disk and introduce two new wake-up strategies. Our main result is an improved makespan, reducing the previous bound from $7.07$ to $5.4162$, as established in Bonichon et al. (2024a). Formally, we state the following theorem:
@@
A robot at the origin can wake up any set of $n$ asleep robots in the unit $l_{2}$-disk with a makespan of at most $5.4162$. The proof of this theorem is provided in Section 3.1. Also, concurrently with our work, Bonichon et al. (2024b) reported a ratio of $4.63r$ for $(\mathbb{R}^{2},l_{2})$. Our first strategy is similar to their strategy; however, our second strategy is completely different. Next, we examine FTP in $\mathbb{R}^{3}$ and propose a new strategy for $(\mathbb{R}^{3},l_{1})$. We establish an upper bound of $13$ for $(\mathbb{R}^{3},l_{1})$, which leads to an upper bound of $13\sqrt{3}$ for $(\mathbb{R}^{3},l_{2})$. To our knowledge, no previous bounds for the makespan in $\mathbb{R}^{3}$ have been provided. Formally, we state the following theorem:
@@
A robot at the origin can wake up any set of $n$ asleep robots in the unit $l_{1}$-ball in $\mathbb{R}^{3}$ with a makespan of at most $13$. The proof of this theorem is presented in Section 4. Our approaches for $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ are fundamentally different. In $\mathbb{R}^{2}$, we start at the origin and initially awaken robots within a disk of radius zero, since only the origin robot is awake. In each step, we expand the radius of this disk, waking up the robots that are closer to the origin, until the radius reaches $1$. By the end, all robots inside the disk are awake. However, this method cannot be extended to $\mathbb{R}^{3}$. Therefore, we employ a different strategy for $\mathbb{R}^{3}$.
@@
Finally, we examine a version of FTP in $(\mathbb{R}^{3},l_{2})$, where the asleep robots are located on the boundary of the unit $l_{2}$-ball. Using our approach for $(\mathbb{R}^{2},l_{2})$, we show that FTP in $(\mathbb{R}^{3},l_{2})$ for asleep robots on the boundary can be solved with a makespan of at most $12.37$. Formally, we have: {theorem} A robot at the origin can wake up any set of $n$ asleep robots on the boundary of the unit $l_{2}$-ball in $\mathbb{R}^{3}$ with a makespan of at most $12.37$. The proof of this theorem is deferred to Section 5.2. Our approach involves mapping the points on the boundary of the unit $l_{2}$-ball in $\mathbb{R}^{3}$ to a $l_{2}$-disk in $\mathbb{R}^{2}$. By solving the problem in $\mathbb{R}^{2}$, we obtain a wake-up strategy for $\mathbb{R}^{3}$. One motivation for tackling the problem in this setting is our conjecture that the maximum makespan for $n$ points is reached when the points lie on the boundary of the unit $\eta$-ball.
A similar wake-up strategy allows us to tackle another variant of FTP, which we call surface-FTP. In this version, the points are positioned on the boundary of a unit $l_{2}$-ball in $\mathbb{R}^{3}$, with the active robot also on the surface. The distance between two points $v$ and $u$ is defined as the geodesic distance, or the length of the shortest arc connecting them on the surface. This problem has practical applications in areas like communication and transportation on Earth’s surface, where geodesic distance between points is important. Formally, we have: {theorem} Given an instance of surface-FTP, the makespan is at most $11.65$. This theorem is proven in Section 5.3.
@@ 3. FTP in $(\mathbb{R}^{2},l_{2})$
3.1. The wake-up ratio is at most $5.4162$ in $(\mathbb{R}^{2},l_{2})$
This section establishes an upper bound of $5.4162$ for $\gamma_{({\mathbb{R}^{2}},l_{2})}$. We present two key strategies for computing a wake-up tree and combine them to achieve an improved makespan.
@@ 3.1.1. Arc-Strategy
The Arc-Strategy provides an upper bound of $7.9651$ for the wake-up ratio in $(\mathbb{R}^{2},l_{2})$.
@@
$\sum_{i=1}^{k}{d_{i}}=d_{1}+d_{2}+d_{3}+\sum_{i=4}^{k}{\sqrt{b_{i}^{2}+a_{i}^{2}+2a_{i}b_{i}\sin{\frac{\alpha_{i-1}}{2}}}}$
@@
$\sum_{i=1}^{k}{d_{i}}$ $=d_{1}+d_{2}+d_{3}+$ $\sum_{i=4}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}(\sin{\frac{\alpha_{i-1}}{2}})^{2}}}$
@@
$\sum_{i=4}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}(\sin{\frac{\alpha_{i-1}}{2}})^{2}}}$
$\leq\sum_{i=4}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}}$
$=\sum_{i=4}^{k}{\sqrt{(a_{i}+2r_{i-1}\sin{\frac{\alpha_{i-1}}{2}})^{2}}}=\sum_{i=4}^{k}{a_{i}+2r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}$
@@
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $d_{1}+d_{2}+d_{3}+\sum_{i=4}^{k}{a_{i}}+2\sum_{i=4}^{k}{r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}$
Note that $\sum_{i=4}^{k}a_{i}=1-r_{3}$ and $d_{1}=r_{1}$ and $d_{2}\leq r_{1}+r_{2}$ and $d_{3}\leq r_{2}+r_{3}$ using triangle inequality so by replacing them and considering all $r_{i}\leq 1$ we have:
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $r_{1}+r_{1}+r_{2}+r_{2}+r_{3}+1-r_{3}+2\sum_{i=3}^{k}{\sin{\frac{\alpha_{i}}{2}}}$
Given that $\alpha_{i}\leq\frac{\pi}{2^{i-2}}$ for $i>1$, we can write:
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2r_{1}+2r_{2}+2\left(\sin{\frac{\pi}{4}}+\sin{\frac{\pi}{8}}+\dots\right)$
Using the fact that for $x\geq 0$, $\sin{x}\leq x$, we have:
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2r_{1}+2r_{2}+2\left(\sin{\frac{\pi}{4}}+\sin{\frac{\pi}{8}}\right)+2\pi\left(\frac{1}{16}+\dots+\frac{1}{2^{n}}\right)$
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\left(\frac{\sqrt{2}}{2}+0.3827\right)+2\pi\left(\frac{1}{16}+\dots\right)+2r_{1}+2r_{2}$
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+\left(0.7654+1.4143\right)+\frac{\pi}{4}+2r_{1}+2r_{2}$
$\sum_{i=1}^{k}{d_{i}}$ $<$ $3.9651+2r_{1}+2r_{2}\leq 7.9651$
[…]
A figure of a path in the Arc-Strategy. This path starts from the center and visits the points in the order specified by the arc strategy.
@@
Thus, using the Arc-Strategy, FTP in $(\mathbb{R}^{2},l_{2})$ can be solved in at most $7.9651r$ time units, where $r$ is the radius of the disk containing all the asleep robots, and the active robot is at the center.
@@ 3.1.2. Ring-Strategy
  • Consider the projections of the points in $p_{0}$’s half of the ring onto the circle $C(O,r_{1})$. The projection of a point $p_{j}$ onto the circle $C(O,r)$ is the intersection of $\overline{Op_{j}}$ with the circle $C(O,r)$. Let $p_{1}$ be the point whose projection is closest to $p_{0}$. Then, $p_{0}$ activates $p_{1}$.
@@
We now calculate the wake-up time for this strategy, specifically the time required to activate all the robots in a half-ring when the active robot is $p_{0}$. Since there are two awake robots at the same initial position, $p^{\prime}_{0}$ activates the robots in its own half in parallel. Therefore, the upper bound on the wake-up time is the same for both halves, and all robots will be awake within this time.
@@
[…]
The figure shows a path obtained by the Ring-Strategy. The ring is drawn based on closest point to the center of the shape.
@@ 3.1.3. Our combined strategy for FTP in $(\mathbb{R}^{2},l_{2})$
In the first step, $p_{0}$ moves toward $p_{1}$, and now we have two awake robots. Divide the disk into two halves by drawing the diameter that passes through $\overline{Op_{1}}$. Now we have two active robots in the location of $p_{1}$ and each robot is responsible for one half.
Now, assume we want to wake the right half using $p_{0}$. Let $p_{2}$ be the second nearest point to the center in the right half of the ring and $r_{2}$ represent its distance from the center. If $r_{2}\leq 0.3627$, then using the Arc-Strategy, this half can be activated in $2r_{1}+2r_{2}+3.9651\leq 4r_{2}+3.9651\leq 5.4162$ time units, see the proof of Lemma 3.1.1. Otherwise, using the Ring-Strategy, $p_{0}$ moves $r_{2}-r_{1}$ units in the direction of $\overrightarrow{Op_{1}}$. The half-ring on the right now has an inner radius of $r_{2}$ and an outer radius of $1$. So, we can activate the robots in this half-ring in $r_{1}+(r_{2}-r_{1})+3+\pi-3r_{2}<6.1416-2\times 0.3627=5.4162$ time units, see the proof of Lemma 3.1.2.
Similarly, let $p_{2^{\prime}}$ be the second nearest point to the center in the left half of the disk. Again, based on the length of $r_{2^{\prime}}$, we choose the best strategy for activating the robots in this half, which takes at most $5.4162$ time units. Thus, the proof is complete.
@@ 3.2. Some examples
We show that for certain values of $n$, the wake-up ratio for $n$ points is not achieved when the points are equally distributed on the unit circle by presenting specific instances and computing their makespan. Our first instance is a set of $5$ asleep robots, the makespan of this instance is $3.530$, see Figure 5. When five points are equally distributed on the cycle then the makespan is $3.351$. Our second example is a set of $7$ asleep robots, the makespan of this instance is $3.498$, see Figure 5. When seven points are equally distributed on the cycle then the makespan is $3.431$. The makespans of these instances are computed using a computer program by brute force search. Our code is publicly available online [^1] Makespan Calculator Code.
@@
Point $\theta$ x y
$p_{1}$ $17.37$ $0.954$ $0.298$
$p_{2}$ $96.22$ $-0.108$ $0.994$
$p_{3}$ $174.42$ $-0.995$ $0.097$
$p_{4}$ $219.75$ $-0.768$ $-0.639$
$p_{5}$ $299.27$ $0.488$ $-0.872$
@@
Furthermore, we generated over $10^{5}$ random instances for different values of $n$ and computed their makespans, leading to the following conjecture: